OpenStudy (bananas1234):

What is the area of a circle whose radius is 10 meters? 20 pi m^2 25 pi m^2 100 pi m^2 400 pi m^2

OpenStudy (bananas1234):

@Ahsome

OpenStudy (ahsome):

What is the equation for the area of a circle?

OpenStudy (bananas1234):

pi r^2

OpenStudy (sepeario):

if the equation for the area as you stated is pi r ^2. then, substitute the value of the radius: 10. therefore (pi 10 ^ 2)^2

OpenStudy (sepeario):

hope i helped!

OpenStudy (bananas1234):

C?

OpenStudy (sepeario):

yes.

OpenStudy (bananas1234):

awesome! thx!!

OpenStudy (ahsome):

\[\Huge{Area=\pi\times r^2}\]\[\Huge{Area=\pi\times 10^2}\]\[\Huge{Area=\pi\times 100}\]\[\Huge{Area=100\pi \text{ }m^2}\]Therefore, its C

OpenStudy (bananas1234):

thanks guys! @Ahsome and @Sepeario

OpenStudy (ahsome):

NP

OpenStudy (sepeario):

my pleasure.